Write the inequality , 3/2<x+1<7/2 in the modulus form


1
Expert's answer
2021-04-15T07:34:14-0400

Given that


"\frac{3}{2}<x+1<\frac{7}{2} \newline

\therefore\:\frac{3}{2}-1<x+1-1<\frac{7}{2}-1"


"\frac{1}{2}<x<\frac{5}{2}"


The range of x belong to (1/2 , 5/2).

Now,

"\qquad \frac{1}{2}<x<\frac{5}{2}\newline
\therefore \quad \frac{1}{2}-\frac{3}{2}<x-\frac{3}{2}<\frac{5}{2}-\frac{3}{2} \newline
\therefore \quad -1<x-\frac{3}{2}<1\newline
\therefore \quad |x-\frac{3}{2}|<1"


Hence the given ineqaulity "\frac{3}{2}<x+1<\frac{7}{2}" in the modulus form is "|x-\frac{3}{2}|<1"